40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 10:52:38 2019, gautam, Update, SUS, EY door opened 
    Reply  Tue Jan 15 15:49:47 2019, gautam, Update, SUS, EY door opened EY_OSEMs.png
       Reply  Sat Jan 19 15:33:18 2019, gautam, Update, SUS, ETMY OSEMs faulty Screen_Shot_2019-01-19_at_3.32.35_PM.pngETMY_sensors_1231832635.pdf
          Reply  Tue Jan 22 20:36:53 2019, gautam, Update, SUS, ETMY OSEMs faulty 
             Reply  Wed Jan 23 12:39:18 2019, gautam, Update, SUS, EY chamber work 6x
                Reply  Wed Jan 23 23:12:44 2019, gautam, Update, SUS, Prep for FC cleaning IY_level_before.pdf
                   Reply  Thu Jan 24 15:32:31 2019, gautam, Update, SUS, Y arm cavity side first contact applied Yarm_FC.pdfOSEMfilter.png
                      Reply  Tue Jan 29 22:12:40 2019, gautam, Update, SUS, Alignment prep 
                         Reply  Wed Jan 30 11:54:24 2019, gautam, Update, SUS, More alignment prep IMG_5958.JPGIMG_5962.JPG
                            Reply  Wed Jan 30 19:25:40 2019, gautam, Update, SUS, Xarm cavity alignment Xarm.pngXarm_IR.png
                               Reply  Fri Feb 1 01:24:06 2019, gautam, Update, SUS, Almost ready for pumpdown tomorrow 
                                  Reply  Fri Feb 1 13:16:50 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                     Reply  Fri Feb 1 21:52:57 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                  Reply  Fri Feb 1 14:44:14 2019, gautam, Update, SUS, Y arm FC cleaning and reinstall IMG_5974.JPGIMG_5986.JPGIMG_5992.JPG
    Reply  Wed Jan 16 16:25:25 2019, gautam, Update, SUS, Yarm locked Yarm_locked.png
       Reply  Fri Jan 18 21:34:18 2019, gautam, Update, SUS, Unused optic on EY table IMG_5408.JPG
          Reply  Sat Jan 19 05:07:45 2019, Koji, Update, SUS, Unused optic on EY table 
Message ID: 14411     Entry time: Tue Jan 22 20:36:53 2019     In reply to: 14409     Reply to this: 14413
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY OSEMs faulty 

Short update on latest Satellite box woes.

  1. I checked the resistance of all 5 OSEM coils on ETMY using a DB25 breakout board and a multimeter - all were between 16-17 ohms (mesured from the cable to the Vacuum flange), which I think is consistent with the expected value.
  2. Checked the bias voltage (aka slow path) from the coil driver board was reaching the coils
    • The voltages were indeed being sent out of the coil driver board - I confirmed by driving a slow sine wave and measuring at the output of the coil driver board, with all the fast outputs disabled.
    • The voltage is arriving at the 64 pin IDC connector at the Satellite box - Chub and I verified this using some mini-grabbers and leads from wirewound resistors (we don't have a breakout board for this kind of connector, would be handy to get some!)
    • However, the voltages are not being sent out through the DB25 connectors on the side of the Satellite box, at least for the LL and UR channels. UL seems to work okay.
    • This behavior is consistent with the observation that we had to apply way larger bias voltages to get the cavity axis to line up than was the nominal values - if one or more coils weren't getting their signals, it would also explain the large PIT->YAW coupling I observed using the Oplev spot and the slow bias alignment EPICS sliders.
    • This behavior is puzzling - the Sat box is just supposed to be a feed-through for the coil driver signals, and we measured resistances between the 64 pin IDC connector and the corresponding DB25 pins, and measured in the range of 0.2-0.3 ohms. However, the voltage fails to make it through - not sure what's going on here.. We will investigate further on the electronics bench.

What's more - I did some Sat box switcheroo, swapping the SRM and ETM boxes back and forth in combination with the tester box. In the process, I seem to have broken the SRM sat box - all the shadow sensors are reporting close to 0 volts, and this was confirmed to be an electronic problem as opposed to some magnet skullduggery using the tester box. Once we get to the bottom of the ETMY sat box, we will look at SRM. This is more or less the last thing to look at for this vent - once we are happy the cavity axis can be recovered reliably, we can freeze the position of the elliptical reflector and begin the F.C.ing.

ELOG V3.1.3-